OpenStudy (anonymous):

Dylan was given a gift card for $100 to an online music store. Singles cost $0.99 and albums cost $9.99. Which of the following is a combination of the number of singles and albums Dylan could have purchased without spending more than the value of the gift card? 10 singles and 11 albums 11 singles and 10 albums 12 singles and 9 albums 13 singles and 8 albums

OpenStudy (anonymous):

singles are .99 and albums are 9.99 (a) 10(.99) + 11(9.99) = 9.90 + 109.89 = 119.79 (too much) (b) 11(.99) + 10(9.99) = 10.89 + 99.90 = 110.79 (too much) all you have to do is just plug in the number of singles and the number of albums and multiply by the cost of the singles and albums and match the one that is $100 or less....you can finish this problem. I have already eliminated two of them for you.

OpenStudy (anonymous):

I got the answer thank you. :)

OpenStudy (anonymous):

your welcome :)